nolegirl24:

Please Help Me!! Evaluate fourth root of 9 multiplied by square root of 9 over the fourth root of 9 to the power of 5. 9 to the power of negative 1 over 2 9 to the power of negative 1 over 4 9 9^2

Ultrilliam:

@Vocaloid

Vocaloid:

\[\frac{ \sqrt[4]{9}*\sqrt{9} }{ (\sqrt[4]{9})^{5} }\]

Vocaloid:

this? :S

nolegirl24:

Yes! How did you do that? @Vocaloid

Ultrilliam:

Using LaTeX, you can see a tutorial I made here: WARNING IT DOES TAKE A WHILE TO LOAD questioncove.com/updates/5654ccade4b0af360a7ccd28

Vocaloid:

well we can do a bit of re-writing to get:

Vocaloid:

\[\frac{ \sqrt[4]{9}*3 }{ \sqrt[4]{9}*\sqrt[4]{9}*\sqrt[4]{9}*\sqrt[4]{9}*\sqrt[4]{9} }\]

Vocaloid:

the 4th root of 9 times itself 4 times is just 9

Vocaloid:

\[\frac{ \sqrt[4]{9} * 3}{ 9 * \sqrt[4]{9} }\]

Vocaloid:

then we just cross out the last 4th root of 9 and then we just get 3/9 = 1/3

nolegirl24:

But that is not an answer :(

Vocaloid:

oh, well we can re-write 1/3 as 1/sqrt(9) which is just 9^(-1/2)

Vocaloid:

so the first option

nolegirl24:

Oh that makes sense thank you @Vocaloid